Stock in Thanksgiving Industries has a beta of 1.2. The market risk premium is 5 percent, and T-bills are currently yielding 4.9 percent. The company’s most recent dividend was $1.90 per share, and dividends are expected to grow at an annual rate of 5 percent indefinitely.

If the stock sells for $50 per share, what is your best estimate of the company’s cost of equity? Hint: use both approaches! (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

cost of equity as per growth model=(D1/Current price)+Growth rate

=[(1.9*1.05)/50]+0.05

=8.99%

cost of equity as per CAPM=risk free rate+Beta*market risk premium

=4.9+(1.2*5)=10.9%

Best estimate of cost of equity=(8.99+10.9)/2

=9.95%(Approx).